A Complete Bibliography of Publications in Mathematics and Computers in Simulation: 1950–1959

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.00

Title word cross-reference

2^{nd} [13].

1958 [1].

25 [55].

Activité [20, 31, 8, 40, 50, 60]. advertising [11]. aerodynamic [53]. aerodynamics [52]. aérodynamique [52]. aléatoire [36]. amplitude [44]. Analog [15, 13, 12, 16, 35, 54, 45, 56, 3, 46, 22, 14, 36]. analogical [23]. Analogie [52, 3]. Analogie-Rechenmaschinen [3]. analogique [54, 15, 12, 16, 46, 22, 35, 36, 10]. analogy [52, 53]. Analyse [25]. analyser [4].

Analysis [25]. any [16]. Application [54, 16, 46, 15]. association [20, 31, 40, 50, 60, 10, 8]. automatic [26]. automatique [26].

band [16]. band-pass [16]. bande [16]. basse [24]. between [26]. bibliographique [6, 18, 29, 38, 48, 58]. broad [46]. broad-spectrum [46].

C.E.R.N. [55]. Calcul [10, 15, 12, 16, 46]. calculateur [36]. calculation [15, 3, 16, 46, 12]. calculator [36]. Chronique [6, 18, 29, 38, 48, 58]. circuits [45]. comparators [44]. Computation [13, 56, 14]. computer [45]. concept [23]. conception [22]. Conditions [11]. correction [44]. d’articles [7, 19, 30, 49, 59]. Days [12]. d’énergie [35]. design [22]. Development [36]. Device [24]. Digital [56]. d’impulsions [46]. Dispositif [24]. Documents [9, 21, 32, 41, 51, 61]. d’un [24, 16, 35].

Editorial [33, 34, 2, 42, 43]. electronic [45, 3]. Elektronische [3]. elektronischen [45]. elevated [24]. élevée [24]. energy [35]. entre [26]. Errata [27]. étude [54, 46, 35]. expansion [15]. expérimentales [26]. expressive [26].

filters [16]. filtres [16]. fonctions [25]. forme [24]. fr [39]. French [11, 9, 21, 32, 26, 41, 51, 61, 24, 54, 15, 25, 12, 16, 46, 52, 22, 35, 36]. fréquence [24]. frequency [24, 55]. function [55]. functions [25].

General [11]. générales [11]. générateurs [46]. génération [36]. generator [55]. generators [45]. German [45, 3]. GeV [55].

hyperstatic [15]. hyperstatiques [15]. Hysteresis [44].

Informations [5, 17, 28, 47, 57, 37]. International [13, 12]. internationale [10]. internationales [12]. iteration [14].

journées [12]. June [1].

kurzschlüsse [45].

large [46]. length [36]. longueur [36]. low [24].

machines [3, 4]. mathematics [26]. mathématiques [26]. Meeting [13]. Method [14, 54, 16]. Méthode [16, 54].

Nachbildung [45]. network [4].

Optimal [3]. Optimales [3].

Pages [1]. pass [16]. passe [16]. passe-bande [16]. phénomène [24]. phenomenon [24]. pipes [15]. plus [24]. pole [4]. pour [10]. procédés [46]. procedures [46]. processes [54]. processus [54]. producing [24]. produisant [24]. Programmieren [3]. programming [55, 3]. propos [12]. proton [55]. publicité [11]. pulsation [16]. pulse [16]. pulsers [46].

quelconque [16].

random [36]. Rapports [26]. Received [9, 21, 32, 41, 51, 61]. Rechenmaschinen [3]. rechner [45]. rectangulaires [36]. rectangular [36]. recus [9, 21, 32, 41, 51, 61]. Relations [26]. repetitive [54, 54]. réponse [16]. reproducing [24]. Reproduction [45]. reproduisant [24]. response [16]. Résumés [7, 19, 30, 49, 59]. Rheoelectric [53]. rhéoélectrique [52]. Rheological [52].

salient [4]. salient-pole [4]. se [24]. Second [12]. secondes [12]. shape [24]. short [45]. signal [16]. signals [36]. signaux [36]. Simulation [4]. sinusoidal [16, 16]. soumis [16]. special [55]. spectre [46]. spectrum [46]. stabilité [35]. stability [35]. stochastic [54]. stochastiques [54]. structures [15]. study [54, 46, 35]. subjected [16]. symétriques [16]. symmetric [16]. synchrongeneratoren [45]. synchronous [45, 4]. synchrotron [55]. synthèse [25]. synthesis [25]. system [55].

thermal [15]. thermique [15]. transfer [25]. transfert [25]. transitoire [16]. transitory [16]. transport [35]. tuyauteries [15]. unsymmetrischer [45]. unsymmmetric [45].

versus [56]. Vitalité [22]. Vitality [23, 22].

References

Anonymous:1958:PJ

[1] Anonymous. Pages 2–56 (June 1958). Mathematics and Computers in Simulation, 1(1):??, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Hoffmann:1958:E

[2] J. Hoffmann. Editorial. Mathematics and Computers in Simulation, 1 (1):2–3, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800014.

Herrmann:1958:OPE

[3] Horst Herrmann. Optimales Programmieren Fur Elektronische Analogie-Rechenmaschinen. (German) [Optimal programming for electronic analog calculation machines]. Mathematics and Computers in Simulation, 1 (1):4–23, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800026.

Vuckovic:1958:SSP

[4] Vladan Vučković and Petar Miljanić. Simulation of salient-pole synchronous machines on the network analyser. Mathematics and Computers in Simulation, 1(1):24–30, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800038.

Anonymous:1958:Ia

[5] Anonymous. Informations. Mathematics and Computers in Simulation, 1 (1):31–38, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S037847545880004X.

Anonymous:1958:CBa

[6] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 1(1):39–46, June 1958. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800051.

Anonymous:1958:RAa

[7] Anonymous. Résumés d’articles. Mathematics and Computers in Simulation, 1(1):47, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800063.

Anonymous:1958:AL

[8] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 1(1):48–54, June 1958. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800075.

Anonymous:1958:DRa

[9] Anonymous. Documents reçus. (French) [received documents]. Mathematics and Computers in Simulation, 1 (1):55, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800087.

Anonymous:1958:AIP

[10] Anonymous. Association internationale pour le Calcul analogique. Mathematics and Computers in Simulation, 1(1):56, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800099.

Anonymous:1958:CGP

[11] Anonymous. Conditions générales de la publicité. (French) [General advertising conditions]. Mathematics and Computers in Simulation, 1(1): CO3, June 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800105.

Hoffmann:1958:PSJ

[12] J. Hoffmann. À propos des secondes journées internationales de calcul analogique. (French) [About the Second International Days of Analog Calculation]. Mathematics and Computers in Simulation, 1(2):58, September 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800221.

Hoffmann:1958:AIA

[13] J. Hoffmann. About the 2^{nd} International Analog Computation Meeting. Mathematics and Computers in Simulation, 1 (2):59, September 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800233.

Tomovic:1958:MIA

[14] Rajko Tomovic. Method of iteration and analog computation. Mathematics and Computers in Simulation, 1(2): 60–63, September 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800245.

Goethals:1958:CAS

[15] J. Goethals. Calcul analogique des structures hyperstatiques: Application au calcul des tuyauteries en expansion thermique. (French) [analog calculation of hyperstatic structures: Application to the calculation of thermal expansion pipes]. Mathematics and Computers in Simulation, 1(2):64– 74, September 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800257.

Jespers:1958:MCA

[16] Paul Jespers. Méthode de calcul analogique de la réponse transitoire des filtres passe-bande symétriques soumis à l’application d’un signal sinusoïdal de pulsation quelconque. (French) [analog calculation method of the transitory response of the symmetric band-pass filters subjected to the application of a sinusoidal signal of any pulse]. Mathematics and Computers in Simulation, 1(2): 75–83, September 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800269.

Anonymous:1958:Ib

[17] Anonymous. Informations. Mathematics and Computers in Simulation, 1(2): 84–85, September 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800270.

Anonymous:1958:CBb

[18] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 1(2):86–101, September 1958. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475458800282.

Anonymous:1958:RAb

[19] Anonymous. Résumés d’articles. Mathematics and Computers in Simulation, 1 (2):102–103, September 1958. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800294.

Anonymous:1958:AAa

[20] Anonymous. Activité de l’association.

Mathematics and Computers in Simulation, 1(2):104–110, September 1958. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475458800300.

Anonymous:1958:DRb

[21] Anonymous. Documents reçus. (French) [received documents]. Mathematics and Computers in Simulation, 1(2):111– 112, September 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800312.

Peretz:1958:VCA

[22] Richard Peretz. Vitalité de la conception analogique. (French) [Vitality of analog design]. Mathematics and Computers in Simulation, 1(3): 114–115, ???? 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800117.

Peretz:1958:VAC

[23] Richard Peretz. Vitality of the analogical concept. Mathematics and Computers in Simulation, 1(3):116– 117, ???? 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800129.

Becar:1958:DRF

[24] Adrien Becar. Dispositif reproduisant à une fréquence basse la forme d’un phénomène se produisant à une fréquence plus élevée. (French) [Device reproducing in a low frequency the shape of a phenomenon producing an elevated frequency]. Mathematics and Computers in Simulation, 1(3):118– 124, ???? 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800130.

Hernalsteen:1958:ASF

[25] P. Hernalsteen. Analyse et synthèse de fonctions de transfert. (French) [Analysis and synthesis of transfer functions]. Mathematics and Computers in Simulation, 1(3):125–131, ???? 1958. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800142.

Anonymous:1958:RAM

[26] Anonymous. Rapports entre l’automatique et les mathématiques expérimentales. (French) [relations between automatic and expressive mathematics]. Mathematics and Computers in Simulation, 1 (3):132–133, ???? 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800154.

Anonymous:1958:E

[27] Anonymous.  Errata.  Mathemat-

ics and Computers in Simulation, 1 (3):134, ???? 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475458800166.

Anonymous:1958:Ic

[28] Anonymous. Informations. Mathematics and Computers in Simulation, 1(3):

135–140, ???? 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800178.

Anonymous:1958:CBc

[29] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 1(3):141–151, ???? 1958. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S037847545880018X.

Anonymous:1958:RAc

[30] Anonymous. Résumés d’articles. Mathematics and Computers in Simulation, 1(3):152, ???? 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800191.

Anonymous:1958:AAb

[31] Anonymous. Activité de l’association. Mathematics and Computers in Simulation, 1(3):153–157, ???? 1958. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475458800208.

Anonymous:1958:DRc

[32] Anonymous. Documents reçus. (French) [received documents]. Mathematics and Computers in Simulation, 1(3): 158–160, ???? 1958. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S037847545880021X.

Gomperts:1959:Ea

[33] R. Gomperts. Editorial. Mathematics and Computers in Simulation, 1(4): 162–163, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800021.

Gomperts:1959:Eb

[34] R. Gomperts. Editorial. Mathematics and Computers in Simulation, 1(4): 164–165, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800033.

Perret:1959:EAS

[35] René Perret. Étude analogique de la stabilité d’un transport d’énergie. (French) [analog study of the stability of an energy transport]. Mathematics and Computers in Simulation, 1(4): 166–180, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800045.

Vichnevetsky:1959:GSR

[36] R. Vichnevetsky. La génération de signaux rectangulaires de longueur aléatoire sur calculateur analogique. (French) [development of rectangular signals of random length on analog calculator]. Mathematics and Computers in Simulation, 1 (4):181–185, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800057.

Schaefer:1959:I

[37] I. A. Schaefer and G. H. Hinds. Informations. Mathematics and Computers in Simulation, 1(4):186–192, ???? 1959. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800069.

Anonymous:1959:CBa

[38] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 1(4):193–202, ???? 1959. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800070.

Anonymous:1959:F

[39] Anonymous.  fr????  Mathemat-

ics and Computers in Simulation, 1 (4):203, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800082.

Anonymous:1959:AAa

[40] Anonymous. Activité de l’association. Mathematics and Computers in Simulation, 1(4):204–205, ???? 1959. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800094.

Anonymous:1959:DRa

[41] Anonymous. Documents reçus. (French) [received documents]. Mathematics and Computers in Simulation, 1(4):

206–217, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800100.

Willems:1959:Ea

[42] Paul Willems. Editorial. Mathematics and Computers in Simulation, 1(5): 220–221, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800112.

Willems:1959:Eb

[43] Paul Willems. Editorial. Mathematics and Computers in Simulation, 1(5): 222–223, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800124.

Tomovic:1959:HCA

[44] Rajko Tomovic and Dusan Mitrovic. Hysteresis correction of amplitude comparators. Mathematics and Computers in Simulation, 1(5):224– 225, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800136.

Hannakam:1959:NUK

[45] Ludwig Hannakam. Nachbildung unsymmetrischer kurzschlüsse von synchrongeneratoren auf dem elektronischen analog rechner. (German) [Reproduction of unsymmmetric short circuits of synchronous generators on the electronic analog computer]. Mathematics and Computers in Simulation, 1(5): 226–242, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800148.

Jespers:1959:APC

[46] Paul Jespers. Une application des procédés de calcul analogique a l’étude des générateurs d’impulsions à large spectre. (French) [an application of the procedures of analog calculation to the study of broad-spectrum pulsers]. Mathematics and Computers in Simulation, 1(5): 243–250, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S037847545980015X.

Anonymous:1959:Ia

[47] Anonymous. Informations. Mathematics and Computers in Simulation, 1(5): 251–254, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800161.

Anonymous:1959:CBb

[48] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 1(5):255–266, ???? 1959. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800173.

Anonymous:1959:RAa

[49] Anonymous. Résumés d’articles. Mathematics and Computers in Simulation, 1(5):267, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800185.

Anonymous:1959:AAb

[50] Anonymous. Activité de l’association.

Mathematics and Computers in Simulation, 1(5):268–269, ???? 1959. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800197.

Anonymous:1959:DRb

[51] Anonymous. Documents reçus. (French)

[received documents]. Mathematics and Computers in Simulation, 1(5): 270–274, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800203.

Malavard:1959:ARA

[52] L. Malavard. Analogie rhéoélectrique et aérodynamique. (French) [Rheological analogy and aerodynamics]. Mathematics and Computers in Simulation, 1 (6):276–277, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800215.

Malavard:1959:RAA

[53] L. Malavard. Rheoelectric and aerodynamic analogy. Mathematics and Computers in Simulation, 1(6):278– 279, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800227.

Gabillard:1959:AMR

[54] Robert Gabillard. Application de la méthode répétitive a l’étude analogique de processus stochastiques. (French) [application of the repetitive method to the analog study of stochastic processes]. Mathematics and Computers in Simulation, 1(6):280–289, ???? 1959. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800239.

Germain:1959:SFG

[55] P. Germain, C. Schmelzer, W. Schnell, and A. Susini. A special function generator: The frequency programming system for the 25 GeV C.E.R.N. proton synchrotron. Mathematics and Computers in Simulation, 1(6):290– 297, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800240.

Harder:1959:DVA

[56] Edwin L. Harder. Digital versus analog computation. Mathematics and Computers in Simulation, 1(6):298– 300, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0378475459800252.

Anonymous:1959:Ib

[57] Anonymous. Informations. Mathematics and Computers in Simulation, 1(6):

301–308, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800264.

Anonymous:1959:CBc

[58] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 1(6):309–320, ???? 1959. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800276.

Anonymous:1959:RAb

[59] Anonymous. Résumés d’articles. Mathematics and Computers in Simulation, 1 (6):321–322, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800288.

Anonymous:1959:AAc

[60] Anonymous. Activité de l’association.

Mathematics and Computers in Simulation, 1(6):323–324, ???? 1959. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S037847545980029X.

Anonymous:1959:DRc

[61] Anonymous. Documents reçus. (French)

[received documents]. Mathematics and Computers in Simulation, 1(6): 325–332, ???? 1959. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0378475459800306.